2. Find the Thevenin equivalent of the circuit shown below and determine the maximum power that it could transfer to a load rPlease show all work thank you
Answer #1

N 2 Vo 12 _ h No a V3: at V : -O N, - V2 = -12 at v tva: V - 2 vo + v + y2 = 30 + v2 [ ++ 4] - v3[+ +2] -0. O at Us: V₂ -V2 -Vo = Utha V₂ = 16V To find Rth short at vo a find IS N Vo=o at loop i 1 - 12 +2(4, -isne) to 34 - Right 12 -0 « Evg Сам? -24arbuit: The vening equivalent . . Rona 3.33 Me E le : 3.333 16V 163 4x3.333. Mont Pmase = 19.20